

********************************************************************
UARS WEEKLY STATUS REPORT - 02 Nov 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all
available orbital tracking periods and stellar calibration observations
whenever possible. We continue to gather enough data to meet the
SOLSTICE science objectives during this period of reduced tape recorder
operations. Furthermore, we have validated the PACOR-A data transfer
files.
***********
SUSIM
***********
SUSIM completed daily mid resolution and MgII solar scans
during periods of predicted UARS telemetry.
SUSIM's CDHF reprocessing remains about 200 days short of
completion.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
***** ALL PEM COMMUNICATION SHOULD NOW BE THROUGH THE PEM2K:: NODE *****
PEM would like to thank the project for their work on convincing NASA HQ
to keep UARS alive and operating in campaign mode.
The PEM boom instruments were activated on 2001303 at about 22 UT in
full operational mode.
Tests of PEM Level 0 data transfer to PEM2K:: from PACOR-A succeed;
however, the file naming convention from PACOR-A is not known. PEM was
expecting to receive files using the same naming convention as is used on
the PEM Level 0 data files received from the CDHF. If anyone knows the
meaning of the PACOR-A file name convention, please inform Sandee Jeffers
at sjeffers@swri.org.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 20 October 2001 (Orbit 55261, GMT Day 293) through 26
October 2001 (Orbit 55366, GMT Day 299).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 293/1317 55269 Power ON (Day-Only)
HALOE 295/0124 55292 Power ON
WINDII 297/0811 55326 Power ON
The beta angle ranged between the angles of 11.2 degrees to 36.8 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 73 degrees on 7 November 2001 (DOY 311).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.9 to 1.0 degrees C. Additional battery performance data for each
battery on 22 October 2001 and 26 October 2001 are provided in Appendix D.
The clock error ranged from Š15.6 to 11.2 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
„ OUT-OF-LIMITS = 00
„ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
„ FOT = 17
„ HALOE = 01
„ WINDII = 01
„ PEM = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
„ FOT = 05
One (1) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 01/267 - 2287.496874
· Transponder B: 01/263 - 2287.498962
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/277 show the observatory data loss to be 303 hours, 18
minutes, and 28.016 seconds. This shows no change in data loss since the last
report period. This is a 0.39 percent data loss, which equals a 99.61 percent
data capture for the mission.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT has begun the process of implementing RaytheonÕs Eclipse ground system
and continues to work on the TAC replacement system. These new systems will
provide enhanced capability and make up for the loss of some UARS facilities,
which have been shut down due to budgetary reasons. This will be a huge effort
and must be completed as soon as possible in order to provide a low cost method
for flying UARS. The FOT will need assistance from all entities involved to
insure the success of this operation. The estimated time for completion is
April 2002.
MEETINGS
The UARS FOT attended an Eclipse-Raytheon All Day Meeting on Tuesday, October
23, 2001.
The UARS FOT attended a UARS Operational Plan Meeting on Wednesday, October 24,
2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 26 November 2001 (DOY 330). This
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 26 October 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
55269 293/1317 01-637 PEM Power ON (Day-Only)
55283 294/1135 01-638 OBC Clk Adj -20.6 to 0.00 msec/day
55288 294/1932 01-639 NBTR NOOPS D-295
55292 295/0124 01-640 HALOE Power ON
55297 295/0953 01-639 OBC NOOPS D-295
55302 295/1749 01-641 NBTR NOOPS D-296
55311 296/0810 01-641 NBTR NOOPS D-296
55319 296/2114 01-641 NBTR NOOPS D-296
55320 296/2254 01-642 NBTR NOOPS D-297
55326 297/0811 01-642 NBTR NOOPS D-297
55326 297/0811 01-643 WINDII Power ON
55332 297/1645 01-644 OBC Clk Adj 0.00 to 20.6 msec/day
55332 297/1750 01-645 NBTR NOOPS D-298
55341 298/0811 01-645 NBTR NOOPS D-298
55349 298/2113 01-646 NBTR NOOPS D-298
55352 299/0035 01-646 NBTR NBTR PBK
55352 299/0122 01-647 OBC Clk Adj 20.6 to 0.00 msec/day
55355 299/0633 01-646 NBTR NOOPS D-299
55363 299/1930 01-648 NBTR NOOPS D-300
55364 299/2113 01-648 NBTR NOOPS D-300
__________________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 26 October 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 26 October 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
55283 294/1140 Clk Adj -20.6 to 0.00 msec/day
55332 297/1647 Clk Adj 0.00 to 20.6 msec/day
55352 299/0125 Clk Adj 20.6 to 0.00 msec/day
_____________________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 26 October 2001
22 October 2001 - GMT Day 295 Beta = 21.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.57/-4.94 +0.0/ -0.0 11.3 0.000 0.0
2 +00.0/ -22.4 +5.15/+3.55 +28.8/-14.5 29.1 1.040 20.9
3 +22.4/ -50.4 +4.19/+2.60 +28.1/-14.3 29.0 1.030 20.5
26 October 2001 - GMT Day 299 Beta = 36.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.57/-4.94 +0.0/ -0.0 11.3 0.000 0.0
2 +00.0/-134.4 +5.46/+4.50 +27.2/-14.4 26.8 1.020 20.7
3 +22.4/ -56.0 +4.50/+3.55 +26.8/-14.8 26.4 1.020 20.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 26 October 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-061 FOT 55268 293/1135 14170 Generic Late Acquisition
01-062 FOT 55287 294/1719 24085 Incorrect Data Time-tag
01-063 FOT 55314 296/1240 N/A TAC Crashed
01-064 FOT 55319 296/2114 N/A Mis-configure
01-065 FOT 55335 297/2254 N/A Unable to Command
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-062 FOT 55287 294/1719 24085 Incorrect Data Time-tag
01-063 FOT 55314 296/1240 N/A TAC Crashed
01-064 FOT 55319 296/2114 N/A Mis-configure
01-065 FOT 55335 297/2254 N/A Unable to Command
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- --- --- -----------------
01-061 FOT see list Oct 01 14170 Generic Late Acquisition
SEVNTFW01/43